Circular No. 8002

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2002 T6 (NEAT-LINEAR) An object reported as asteroidal independently by the NEAT and LIN-

EAR surveys, and later placed on the NEOCP, has been found cometary in appearance. P. Birtwhistle, Great Shefford, U.K., reports that CCD images taken with a 0.3-m reflector on Oct. 27.9 UT show two faint, thin, straight tails in p.a. 107^{o} (36′′ long) and 310^{o} (39′′ long) and a nuclear condensation with diameter ∼ 8′′. T. B. Spahr reports that 5-min exposures taken with the Mt. Hopkins 1.2-m reflector on Oct. 29.2 show a small coma of diameter ∼ 6′′ and an extremely faint tail ∼ 10′′ long in p.a. 330^{o}.

2002 UT α_{2000} δ_{2000} m_{2} Observer
Oct. 4.49860 1^{h}15^{m}17.^{s}65 +10^{o}51′00′′.9 20.1 NEAT
7.25970 1 14 01.38 +10 36 52.1 19.4 LINEAR

The available astrometry, elliptical orbital elements given below, and an ephemeris appear on MPEC 2002-U42.

T = 2003 June 27.8073 TT  ω = 217.^{o}6554
e = 0.556711  Ω = 209.0260 }2000.0
q = 3.390064 AU  i = 11.0366
a = 7.647520 AU n^{o} = 0.0466040 P = 21.15 years

SUPERNOVA 2002he IN UGC 4322 Further to IAUC 7999, J. Graham and W. D. Li report the LOTOSS

discovery of an apparent supernova on unfiltered KAIT CCD images taken on Oct. 28.5 (mag ∼ 17.5) and 29.5 UT (mag ∼ 17.1). SN 2002he is located at α = 8^{h}19^{m}58^{s}.83, δ = +62^{o}49′13′′.2 (equinox 2000.0), which is 19′′.7 west and 37′′.3 south of the nucleus of UGC 4322. A KAIT image taken on Oct. 22.5 showed nothing at this position (limiting mag ∼ 19.0).

SUPERNOVAE 2002hb AND 2002hd M. Hamuy, Carnegie Observatories, reports that a spectrum (range

360–900 nm) of SN 2002hb (cf. IAUC 7999), obtained on Oct. 29.13 UT with the Las Campanas Baade 6.5-m telescope (+ LDSS2), shows it to be a type-Ia supernova, 3–4 weeks past maximum light, at a redshift of ≈ 0.09. A spectrum of SN 2002hd, obtained with the same instrument on Oct. 29.35, reveals it to be a type-Ia supernova, ∼ 5 days past maximum light, at a redshift of ≈ 0.035.

2002 October 29  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green